OpenStudy (anonymous):

The volume of two similar solids is 1331 m3 and 729 m3. The surface area of the larger solid is 605 m2. What is the surface area, in square meters, of the smaller solid?

OpenStudy (anonymous):

A)121 B)305 C)405 D)81

OpenStudy (rajee_sam):

Volume of B (Larger Solid) = (Scale Factor)^3 . Volume of A(Smaller Solid) From this you can find Scale Factor ( Scale Factor is the factor by which the B is larger than A) Surface Area of B = (Scale Factor)^2 . Surface Are of A Please go ahead and solve the problem.

OpenStudy (anonymous):

@rajee_sam can you please dumb it down im not understanding

OpenStudy (rajee_sam):

when you have two similar solids one is bigger or smaller than the other in dimensions. Do you agree?

OpenStudy (anonymous):

yes

OpenStudy (rajee_sam):

the factor that it is larger are smaller is called the scale factor

OpenStudy (rajee_sam):

the dimension of each side I mean

OpenStudy (anonymous):

ok

OpenStudy (rajee_sam):

Now if you consider a square that has its side as 2 and another square that has its side a 6 the scale factor is 3 because 2 x 3 = 6

OpenStudy (rajee_sam):

Now if you find the area of the smaller square it will be 2 x 2 = 4 That of the larger square is 6 x 6 = 36 Now if you compare the areas larger one is 9 times bigger than the smaller which is (scale factor )^2 = 3^2 = 9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hey Isabella, are you ignoring my messages?, just wanna knoe

OpenStudy (rajee_sam):

So Area of Solid B = (Scale Factor)^2 x Area of Solid A For Area you multiply 2 dimensions so you have (Scale Factor)^2 Now you have to extend this concept to a 3D shape. the volume is you multiply 3 dimensions. So Volume of Solid B = (Scale Factor)^3 x Volume of A

OpenStudy (anonymous):

okay

OpenStudy (rajee_sam):

now your problem 1331 = (Scale Factor)^3 x 729 \[\frac{ 1331 }{ 729 } = (Scale Factor)^{3}\]

OpenStudy (rajee_sam):

\[(Scale Factor)^{3} = 1.8258\]

OpenStudy (rajee_sam):

\[Scale Factor = \sqrt[3]{1.8258}\]

OpenStudy (anonymous):

1331.0082

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

so S.A of B = 0.00176 ?

OpenStudy (rajee_sam):

605 = Scale factor ^2 x S.A of Smaller Solid

OpenStudy (rajee_sam):

Sorry Cube root of 1.8258 is 1.22

OpenStudy (rajee_sam):

Scale Factor is 1.22

OpenStudy (rajee_sam):

\[\frac{ 605 }{ 1.22 \times 1.22 } = Surface Area of Smaller Solid\]

OpenStudy (anonymous):

S.A of the smaller solid =404.46

OpenStudy (rajee_sam):

yes

OpenStudy (rajee_sam):

406.46

OpenStudy (rajee_sam):

close

OpenStudy (rajee_sam):

Answer Choice is C

OpenStudy (anonymous):

thank you so much :)